OpenStudy (anonymous):

2x squarded - 3x -5 =0

OpenStudy (anonymous):

2x^2-3x-5=0 factor and find the two 0's

OpenStudy (anonymous):

yes that's the question

OpenStudy (whpalmer4):

okay, work this one just like the one we just did: multiply the earpiece numbers, and find a pair of numbers that multiply to that number and add to -3

OpenStudy (anonymous):

if i'm not mistakened it would be 2 multiplied by -5 which would make it -10?

OpenStudy (whpalmer4):

Good so far. Are you saying 2* -5 because those are the numbers on the equation, or because you've already figured out that 2+ -5 = -3?

OpenStudy (anonymous):

yeah for both of em so it would then become 2x squared + 2 - 5 - 5?

OpenStudy (whpalmer4):

right you are, so long as you stick in the x's in the middle \[2x^2 + 2x - 5x - 5\]

OpenStudy (anonymous):

right forgot the x's then parenthesis? that would make it (2x squared + 2x) - (5x -5)

OpenStudy (whpalmer4):

Okay, now here's a trap for the unwary: when you put parentheses around that second group, what you're really doing is writing \[(2x^2+2x) + (-5x-5)\] see the difference?

OpenStudy (anonymous):

but don't you put that on the outside then just change the sign on the inside?

OpenStudy (whpalmer4):

the reason for that is that \(-(5x-5)\) is shorthand for \(-1(5x-5)\) and \[-1(5x-5) = -1*5x -1*-5 = -5x + 5\]

OpenStudy (anonymous):

please explain i am confused

OpenStudy (whpalmer4):

Okay, we don't have any question about the left hand group, it's just the \(-5x-5\) that is causing heartburn, right?

OpenStudy (anonymous):

yup

OpenStudy (whpalmer4):

there are a couple of ways we can deal with this. 1) if you have a choice, put the "middle" numbers in the order that makes this not happen. For example, we could have written\[2x^2 -5x + 2x -5\]and we wouldn't have to worry because we'd just have \[(2x^2-5x) + (2x-5)\] Unfortunately, sometimes you can't do that. 2) just take it for granted that you always want a + between the two groups, and write the ( ) around the whole second thing: \[2x^2 + 2x - 5x -5\]\[(2x^2 + 2x ) + (-5x-5)\] This will always work. The next step is the factoring, and here we'll have \[2x(x+1) -5(x+1)\]right?

OpenStudy (anonymous):

but at first you would have -(5x-1) then you would put that - in the problem and it would become (5x+1) right?

OpenStudy (anonymous):

but we could do it your way and it would come out like that i think then it would become (2x-5) (x+1)

OpenStudy (whpalmer4):

sorry, got a phone call...

OpenStudy (whpalmer4):

so yes, \((2x-5)(x+1)\) is the correct factoring.

OpenStudy (anonymous):

i'm stumped

OpenStudy (whpalmer4):

going back to your mistake earlier: (2x squared + 2x) - (5x -5) \[(2x^2 + 2x) - (5x -5)\]the problem is that the - sign in front of the parentheses means "-1", so what we really have is \[2x^2+2x-1(5x-5)\]and if you use the distributive property to expand that, \[2x^2+2x-1*5x-1(-5) = 2x^2+2x-5x+5 = 2x^2 -3x+5\]and that is not what we started with. We can't do anything while factoring that actually changes the value of the expression.

OpenStudy (anonymous):

oh

OpenStudy (whpalmer4):

That one little sign change transforms it into an expression that cannot be factored, so that's not good!

OpenStudy (whpalmer4):

So, the safest thing (and what I personally do, even after having done hundreds of these problems) is to just always put a + in the middle, and wrap the parentheses around the - signs : \[2x^2 + 2x -5x -5 \]\[(2x^2+2x) + (-5x -5)\]\[2x(x+1) + (-5)(x+1)\]\[2x(x+1)-5(x+1)\]\[(2x-5)(x+1)\] any questions about that?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i'm stumped on the (2x-5) and how you make it smaller where it's the x=

OpenStudy (whpalmer4):

oh, \[(2x-5) (x+1) = 0\]\[(x+1) = 0\]\[(2x-5) = 0\]those last two lines come from the zero product principle, which says that if \(a*b =0\) then either \(a=0\) or \(b=0\) or both. Is your question how to solve \[2x-5=0\]?

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

Okay, put on your helmet, you're going to want to hit your forehead on something hard in a second :-)

OpenStudy (whpalmer4):

\[2x-5=0\]add 5 to both sides \[2x-5+5 = 0+5\]\[2x = 5\]divide both sides by 2 \[\frac{2x}{2} = \frac{5}{2}\]\[x = \frac{5}{2} = 2\frac{1}{2} = 2.5\]

OpenStudy (anonymous):

you've gotta be crapping me

OpenStudy (whpalmer4):

personally, my preference is to keep the first one (5/2) unless you need a decimal number. the middle one is just a PITA

OpenStudy (whpalmer4):

Like I said :-)

OpenStudy (anonymous):

man i'm gonna start calling you doc cause your a life saver

OpenStudy (anonymous):

it would end up being x= 2.5 and x= -1 correct?

OpenStudy (anonymous):

can you help me with some more on here? it's just like four more

OpenStudy (whpalmer4):

let's check them out: \[2x^2-3x-5=0\]\[2(2.5)^2-3(2.5)-5=0\]\[2*6.25 - 7.5-5 = 0\]\[12.5-12.5=0\] \[2(-1)^2-3(-1)-5=0\]\[2+3-5=0\]Yep, good solutions!

OpenStudy (whpalmer4):

I can help you some more, but I also have to make my dinner. Can you take a shot at them and I'll check your answers? Also, if you want to just quickly post all 4, I can have a look and see if there are any traps waiting for you...

OpenStudy (anonymous):

got you x squared = 3x + 40 it would end up being  which would make it have to be multiplied by 40... then you would find a multiple of 40 that adds to 3... erm would that be....i'm stumped